# Improving Speech Emotion Recognition with Mutual Information Regularized Generative Model

### v2

@@ 5.1 SER experiment
Sahu et al. [16] 59.42 60.29
Bao et al. [39] 59.48±0.71 60.37±0.70
Latif et al. [40] 60.51±0.57 61.05±0.68
Malik et al. [18] 58.62±2.11 61.22±1.85
@@
The results depict that our augmentation method achieves the state-of-the-art performance among existing methods. Furthermore, our augmentation method shows highest improvement against its baseline after augmentation, which is $2.6\%$, which is same amount of improvement with Malik et al. [18]. Sahu et al. employed typical emotional label conditional GAN and Bao et al. [39] CycleGAN approach to generate additional data to train SER model. Latif et al. [40] achieved generalization improvement by using unlabelled dataset during training. Still, the improvement of accuracy against their baseline was marginal. Recently, the margin between baseline and data augmented SER was improved when data generation model was conditioned on text input as in Malik et al. [18]. And our model also achieved same margin and state-of-the-art performance by using mutual information between generated data and text data.
@@ 5.2 Multimodal SER Experiment
Multimodal SER experiment was performed on three datasets. This experiment is an use-case-scenario, where you are provided with two emotion prediction models, that one takes speech input and text input for the other. Through this experiment, we intend to show that even with simple data augmentation method via generative model, we improve the performances to the decent amount.
@@ 5.2.1 IEMOCAP
sc_LSTM [41] 55.96 63.51 68.45
bc_LSTM [41] 57.43 67.69 73.89
AE [42] 56.79 25.44 26.25
CRA [43] 56.79 27.96 30.13
MMIN [44] 58.23 67.52 74.72
CIF-MMIN [26] 58.44 69.26 75.65
@@
The experiment result revealed the advantage of our data augmentation method, in terms of accuracy. CRA [43], MMIN [44] and CIF-MMIN [26] exploit the dependency between features from different modality, thus achieving decent performance among other existing works. Inspired from these results, we adopted similar methodology, not to recover missing feature (modality) but to generate new multimodal feature sets.
@@ 5.2.2 MSP-IMPROV
sc_LSTM [41] 40.91 32.01 41.36
bc_LSTM [41] 42.53 57.39 58.32
AE [42] 42.69 27.62 38.73
CRA [43] 38.96 28.37 37.97
MMIN [44] 42.71 56.49 60.98
CIF-MMIN [26] 41.56 58.57 61.72
@@ 5.2.3 MSP-Podcast
Unlike, IEMOCAP and MSP-IMPROV, MSP-Podcast is more naturalistic dataset with much larger number of samples. As there weren’t any existing works that ran experiment with similar protocol, we compared our method with emoDARTS [45], which uses neural architecture search and only audio input. Our method performed better than emoDARTS in terms of variance. Furthermore, compared to emoDARTS, our model can be simpler choice to offer during implementation.
Notable point is, that MSP-Podcast has severe class imbalance problem, thus, we generated data samples for the lesser count of emotional classes to make same numbers of samples per emotion. This scheme of data generation is unlike the generation in existing works, as [29], which requires other modality features to transfer emotional information. In this case, we generated audio feature and text feature from scratch, with only emotional label provided. Still, our data augmentation method synthesized decent sample and we were able to mitigate the class-imbalance problem.
@@ 5.3 Qualitative analysis of generated features
We visualized both generated features from InfoGAN and real audio features for qualitative analysis. The data samples from test dataset of MSP-Podcast was used and same number of generated samples were prepared. Scatter plots of the samples, after transformation via t-sne [46], were produced and analysed. It is clearly observable that samples from different classes are placed on different separable regions (refer to Fig.2).
@@ 6 Conclusion
We have presented a mutual-information-regularised generative framework that effectively addresses data scarcity in speech emotion recognition through principled cross-modal augmentation. Unlike conventional approaches that rely on simplistic label conditioning, our method establishes quantifiable dependencies between generated samples and rich multimodal signals, ensuring the emotional coherence and diversity of synthesised data.
The key insights from our work are threefold. First, mutual information regularisation provides both a training objective and a measurable quality metric, enabling controllable generation of emotionally consistent features. Second, cross-modal transfer from text to audio allows for more nuanced emotion modeling, capturing the complex interplay between linguistic content and vocal expression. Third, our unified framework seamlessly extends to multimodal augmentation, generating paired audio-text samples that significantly enhance multimodal SER performance.

### v3

@@ 5.1 SER experiment
Sahu et al. [sahu2020modeling] 59.42 60.29
Bao et al. [bao2019cyclegan] 59.48±0.71 60.37±0.70
Latif et al. [latif2022self] 60.51±0.57 61.05±0.68
Malik et al. [malik2023preliminary] 58.62±2.11 61.22±1.85
@@
The results depict that our augmentation method achieves the state-of-the-art performance among existing methods. Furthermore, our augmentation method shows highest improvement against its baseline after augmentation, which is $2.6\%$, which is same amount of improvement with Malik et al. [malik2023preliminary]. Sahu et al. employed typical emotional label conditional GAN and Bao et al. [bao2019cyclegan] CycleGAN approach to generate additional data to train SER model. Latif et al. [latif2022self] achieved generalization improvement by using unlabelled dataset during training. Still, the improvement of accuracy against their baseline was marginal. Recently, the margin between baseline and data augmented SER was improved when data generation model was conditioned on text input as in Malik et al. [malik2023preliminary]. Our model also achieved same margin and state-of-the-art performance by using mutual information between generated data and text data. From the perspective of range (from mean minus standard deviation to mean plus standard deviation), we observe that the lower end of the range from our model is higher than others, depicting that our model can consistently outperform existing models.
@@ 5.2 Multimodal SER Experiment
Multimodal SER experiment was performed on three datasets. This experiment is an use-case-scenario, where you are provided with two emotion prediction models, that one takes speech input and text input for the other. Through this experiment, we intend to show that even with simple data augmentation method via generative model, we improve the performances to the decent amount. In this experiment, the previous works [poria2017context, zhao2021missing, liu2024contrastive] did not report the standard deviation, thus, we do not present standard deviation for comparison.
@@ 5.2.1 IEMOCAP
sc_LSTM [poria2017context] 55.96 63.51 68.45
bc_LSTM [poria2017context] 57.43 67.69 73.89
AE [bengio2006greedy] 56.79 25.44 26.25
CRA [tran2017missing] 56.79 27.96 30.13
MMIN [zhao2021missing] 58.23 67.52 74.72
CIF-MMIN [liu2024contrastive] 58.44 69.26 75.65
@@
The experiment result revealed the advantage of our data augmentation method, in terms of accuracy. CRA [tran2017missing], MMIN [zhao2021missing] and CIF-MMIN [liu2024contrastive] exploit the dependency between features from different modality, thus achieving decent performance among other existing works. Inspired from these results, we adopted similar methodology, not to recover missing feature (modality) but to generate new multimodal feature sets. Our text emotion recognition model was rather naive implementation which shows not the best performance, yet, our simple multimodal data augmentation achieved the state-of-the-art.
@@ 5.2.2 MSP-IMPROV
sc_LSTM [poria2017context] 40.91 32.01 41.36
bc_LSTM [poria2017context] 42.53 57.39 58.32
AE [bengio2006greedy] 42.69 27.62 38.73
CRA [tran2017missing] 38.96 28.37 37.97
MMIN [zhao2021missing] 42.71 56.49 60.98
CIF-MMIN [liu2024contrastive] 41.56 58.57 61.72
@@ 5.2.3 MSP-Podcast
Unlike, IEMOCAP and MSP-IMPROV, MSP-Podcast is more naturalistic dataset with much larger number of samples. As there weren’t any existing works that ran experiment with similar protocol, we compared our method with emoDARTS [rajapakshe2024emodarts], which uses neural architecture search and only audio input. Our method performed better than emoDARTS in terms of variance. Furthermore, compared to emoDARTS, our model can be simpler choice to offer during implementation.
Notable point is, that MSP-Podcast has severe class imbalance problem, thus, we generated data samples for the lesser count of emotional classes to make same numbers of samples per emotion. This scheme of data generation is unlike the generation in existing works, as [wang2024incomplete], which requires other modality features to transfer emotional information. In this case, we generated audio feature and text feature from scratch, with only emotional label provided. Still, our data augmentation method synthesized decent sample and we were able to mitigate the class-imbalance problem.
@@ 5.3 Qualitative analysis of generated features
We visualized both generated features from InfoGAN and real audio features for qualitative analysis. The data samples from test dataset of MSP-Podcast was used and same number of generated features were prepared. Scatter plots of the features, after transformation via t-sne [maaten2008visualizing], were produced and analysed. It is clearly observable that features from different classes are placed on different separable regions (refer to Fig.2).
@@ 6 Conclusion
We have presented a mutual-information-regularised generative framework that effectively addresses data scarcity in speech emotion recognition through principled cross-modal augmentation. Unlike conventional approaches that rely on simplistic label conditioning, our method establishes quantifiable dependencies between generated features and rich multimodal signals, ensuring the emotional coherence and diversity of synthesised data.
The key insights from our work are threefold. First, mutual information regularisation provides both a training objective and a measurable quality metric, enabling controllable generation of emotionally consistent features. Second, cross-modal transfer from text to audio allows for more nuanced emotion modeling, capturing the complex interplay between linguistic content and vocal expression. Third, our unified framework seamlessly extends to multimodal augmentation, generating paired audio-text features that significantly enhance multimodal SER performance.